OpenStudy (anonymous):

find x: 3x - 5/6 = 7x+4/4

OpenStudy (anonymous):

This is an equation where similar terms need to be combined. To simplify, look for obvious terms to reduce. Any number divided by itself is 1 so 4/4 = 1. So you now have 3x - 5/6 = 7x +1 Combine like terms, lets keep x positive so move 3x to the Right Hand Side where 7x is and move 1 to the Left Hand Side. The equation is now: -5/6 - 1 = 7x - 3x Finishing the problem: and using a common denominator on the Left Hand Side -5/6 - 6/6 = 4x -11/6 = 4x Divide by 4: -11/24 = x